Progress report: the site in Spring/Summer 2000

January
The first double triple-pendulum suspension (i.e. triple-pendulum suspension for mirror and for reaction mass) has been installed for the near mirror in TCE. The suspension in the east end house has been successfully tested for vacuum suitability.

February
We are starting the installation of the suspension in the power recycling tank. The suspension in TCE has been successfully tested for vacuum suitability. We performed a first test of the data acquisition system by taking data for 8 h from the locked mode cleaner.

March
A horizontal motor has been installed for the length control of mode cleaner 2. A change in control topology has been introduced in order to improve control of the mode-cleaner lengths and of the laser frequency: Instead of a reference cavity the mode cleaners serve as a frequency reference for the laser pre-stabilization. The optimum transmission of the modulation side bands of the power-recycling lock of mode cleaner 2 serves as a dc fequency reference.

April
The new control concept has been tested successfully. The shadow-sensor/coil units of the local controls have been redisigned.

May
The first main mirror (test optics) has been installed in the east end tank TFE. It has been pre-adjusted by means of a He-Ne laser being collinearly reflected through the 600 m arm length.
The active seismic isolation has been demonstrated for the first time in TCE. The horizontal and vertical components of the seismic excitation have been suppressed at the same time by about a factor of ten in the fourier-frequency range of the pendulum resonances.

June
The pockels cells have been replaced by new ones made of LiNbO having a higher modulation index.
During the first half of this year the LabView control programs have been installed. The analog local control electronics has been replaced step by step by LabView controlled ones.

March 2000

Triple pendulum suspension

Suspension for TCE mirror, preassembled on an optical bench. Zoom Image
Suspension for TCE mirror, preassembled on an optical bench.
View from the top: The large cantilever blades. Zoom Image
View from the top: The large cantilever blades.
The suspension has been inserted into tank TCE on top of the stacks. Zoom Image
The suspension has been inserted into tank TCE on top of the stacks.
A look into tank TCE from the top. Zoom Image
A look into tank TCE from the top.
A 3D picture of TCE (to be viewed with cyan-red spectacles). Zoom Image
A 3D picture of TCE (to be viewed with cyan-red spectacles).

Tank TCIb

A test for the suspension of power recycling tank TCIb. Zoom Image
A test for the suspension of power recycling tank TCIb.
View through the RMPR and MPR dummy. Zoom Image
View through the RMPR and MPR dummy.
Stefan adjusting the BDIPR dummy. Zoom Image
Stefan adjusting the BDIPR dummy.

June 2000

Tank TCIb

View from TCC into TCIb. The red spot is the HeNe laser from the east end house. Zoom Image
View from TCC into TCIb. The red spot is the HeNe laser from the east end house.
View into TCIb: BDIPR and reaction mass (l.), MPR and reaction mass (r.), in the background a pick-off mirror for one of the BS beams. Zoom Image
View into TCIb: BDIPR and reaction mass (l.), MPR and reaction mass (r.), in the background a pick-off mirror for one of the BS beams. [less]
View from the top into TCIb. Zoom Image
View from the top into TCIb.

Harald, Stefan and Walter in action

Harald lying in tank TCC during the adjustment of MPR and its reaction mass. Zoom Image
Harald lying in tank TCC during the adjustment of MPR and its reaction mass.
Stefan adjusting BDIPR. Zoom Image
Stefan adjusting BDIPR.
Walter assembling the mode cleaner. Zoom Image
Walter assembling the mode cleaner.

November 2000: Status report

On Nov. 1, the so-called "1200 m cavity" has been locked for the first time. This cavity consists of the power-recycling mirror and one of the interferometer arms. This step took longer than expected, since the windows of the vacuum gate valves caused spatial beam distortions although they had been ordered with very high optical specifications. After pumping down the central cluster and opening the gate valves, the distortions disappeared and the cavity could be locked.Next steps: Over the following two months we will test this cavity and optimize the laser frequency stabilization, measure alignment noise and implement the automatic alignment of the power recycling cavity. Further the electrostatic actuator for the differential armlength control of the Michelson interometer will be tested.

Main Suspensions
The triple pendulum suspensions for the second Michelson arm (the "north arm") will be finished parallel to this, and the installation procedure for the first monolithic suspension will be tested. The pulling machine for the silica fibers has been transferred from Glasgow to Hannover. Geppo has trained some people in Hannover to produce perfect fibers. We expect to install the first monolithic-suspended end mirror in January in the north end building.

Data Acquisition
The data acquisition system is in place and was debugged. It will be used by the scientist on the site to test the installed subsystems. The last missing part in the DAQ chain will be uploaded and tested at the end of November. The part in question is the frame writer that converts the internal GEO file format into the common "FRAME" format used by the other GW projects.

Laser System
The laser is subject to a long-term test in the University lab. No power degradation could be seen for over 60 days. The system will be installed at the site in late December.

December 2000: Status report

The installation of the mode-cleaner system is a success. The mode-cleaner lock works automatically overnight together with the auto-alignment. It runs very regularly with only one reset within fourteen hours. Now we are studying the long-term drift in order to obtain data for a better optimization. The whole system has been stabilized on the power-recycling cavity.

The suspension for the beamsplitter is ready. All the installations in the north-end building are finished, up to the control program (network problems).

Merry Xmas & a Happy New Year to All of the Gravity Wave Community!